# Approximate Polymorphisms of Predicates

### v1

@@ Abstract
We use our techniques to significantly improve the parameter dependence in the work of Friedgut and Regev on approximately intersecting families, from tower type to exponential.
@@ 1 Introduction
The classical BLR linearity test [BLR90] states that if $f\colon\{0,1\}^{n}\to\{0,1\}$ satisfies $\Pr_{x,y\sim\mu_{1/2}}[f(x)\oplus f(y)=f(x\oplus y)]\geq 1-\epsilon$ then there exists a linear function $g\colon\{0,1\}^{n}\to\{0,1\}$ satisfying $\Pr_{x\sim\mu_{1/2}}[g(x)\neq f(x)]\leq\epsilon$; here $\mu_{1/2}$ is the uniform distribution over $\{0,1\}^{n}$. Linearity testing is a special case of a more general problem, approximate polymorphisms, introduced in [CFM+22], which unifies several other related results in the literature, including Kalai’s quantitative Arrow theorem [Kal02] and AND testing [FLMM20].
@@
$\Pr_{\begin{subarray}{c}(x^{(1)}_{1},\dots,x^{(m)}_{1})\sim P\\ \cdots\\ (x^{(1)}_{n},\dots,x^{(m)}_{n})\sim P\end{subarray}}[(f(x^{(1)}),\dots,f(x^{(m)}))\in P]\geq 1-\epsilon.$
@@
In this paper, we extend Theorem 1.2 to arbitrary predicates $P\subseteq\{0,1\}^{m}$ (for all $m$) and to arbitrary distributions $\mu$ on $P$ with full support.
One might hope for a result of the following form: Any approximate polymorphism of $P$ is close to a polymorphism of $P$. While this holds for some predicates $P$, it fails for others, as the following counterexample from [CFM+22] demonstrates.
@@ 1.1 Other results
Theorem 1.5 has an unspecified dependence between $\delta$ and $\epsilon$; the dependence arising from our current proof is of tower type. In two special cases we can improve this dependence.
@@
Let $P\subseteq\{0,1\}^{m}$ be a non-empty monotone predicate: if $x\in P$ and $y\leq x$ (pointwise) then $y\in P$. Let $\mu$ be a distribution on $P$ with full support. There exists a constant $C=C(P,\mu)$ such that the following holds for $\delta=1/\exp\Theta(1/\epsilon^{C})$.
@@
In both cases we get the additional guarantee that if $f_{i}=f_{j}$ and the corresponding marginals of $\mu$ coincide, then $g_{i}=g_{j}$. This implies that if all marginals of $\mu$ are identical then an approximate polymorphism is close to an exact polymorphism.
@@
Fix $20<p<1/$ 2. For every $\epsilon>0$ there exist $\delta>0$ and $j\in\mathbb{N}$ such that the following holds for all $n$ such that $pn$ is an integer.
@@
The dependence of $\delta,j$ on $\epsilon$ is of tower type. We improve the dependence to exponential, at the cost of considering shallow decision trees rather than juntas. (We can convert the decision tree to a junta, losing another exponential in the size of the junta.)
Theorem 1.9 (Improved Friedgut–Regev).
Fix $20<p<1/$ 2. For every $\epsilon>0$ the following holds for all $n$ such that $pn$ is an integer.
If $\mathcal{F}\subseteq\binom{[n]}{pn}$ contains a $1/\exp\Theta(1/\epsilon^{C})$-fraction of the edges of the Kneser graph then there exists an intersecting family $\mathcal{G}\subseteq\binom{[n]}{pn}$ computed by a decision tree of depth $O(1/\epsilon^{C})$ (for some global constant $C$) such that $|\mathcal{F}\setminus\mathcal{G}|\leq\epsilon\binom{n}{pn}$.
@@ 1.2 Related work
Arrow’s celebrated theorem [Arr50] can be expressed in the language of polymorphisms. Let $m\geq 3$. For every permutation $\pi\in S_{m}$, let $I(\pi)\in\{0,1\}^{\binom{m}{2}}$ be the following vector: $I(\pi)(i,j)=[\pi(i)>\pi(j)]$. Arrow’s theorem states that if $f_{i,j}$ is a generalized polymorphism of $P_{m}:=\{I(\pi):\pi\in S_{m}\}$ and each $f_{i,j}$ is unanimous (satisfies $f_{i,j}(b,\dots,b)=b$ for $b\in\{0,1\}$) then there exists $k\in[n]$ such that $f_{i,j}(x)=x_{k}$ for all $i,j$.
Kalai [Kal02] considered the case $m=3$. He showed that if $f_{1,2},f_{2,3},f_{3,1}$ is a $(\mu,\epsilon)$-approximate polymorphism of $P_{3}$, where $\mu$ is the uniform distribution, and furthermore the $f_{i,j}$ are balanced (satisfy $\Pr_{\mu_{1/2}}[f_{i,j}=1]=1/2$), then there exists $k\in[n]$ such that either $f_{i,j}(x)=x_{k}$ for all $i,j$, or $f_{i,j}=1-x_{k}$ for all $i,j$. In this case the predicate $P_{3}$ consists of all triples $(a,b,c)$ such that $a,b,c$ are not all equal. Keller [Kel10] extended Kalai’s result to arbitrary $m\geq 3$ and to arbitrary distributions, under various (severe) restrictions on the $f_{i,j}$.
Mossel [Mos12a] determined all generalized polymorphisms of $P_{m}$ for all $m\geq 3$ (without assuming unanimity), and proved that approximate generalized polymorphisms of these predicates (with respect to the uniform distribution) are close to exact generalized polymorphisms. His techniques in fact work for arbitrary distributions (the missing piece, reverse hypercontractivity for arbitrary distributions, was proved in [MOS13]). Keller [Kel10] improved on Mossel’s result in the case of the uniform distribution by determining the optimal dependence between $\epsilon$ and $\delta$.
@@
Blum, Luby and Rubinfeld [BLR90] were the first to propose the BLR test. They analyzed it using self-correction. Bellare et al. [BCH+96] later gave a different argument using Fourier analysis. The test was generalized to arbitrary prime fields by Kiwi [Kiw03].
David et al. [DDG+17] extended the BLR test to the setting of constant weight inputs, which is analogous to the setting of Theorem 1.9. Translated to the setting of Theorem 1.6, they extended the BLR test to a natural distribution $\mu$ with $\Pr[\mu|_{i}=1]=p$ for all $i$. Dinur et al. [DFH25] analyzed the related “affine test” $f(x\oplus y\oplus z)=f(x)\oplus f(x\oplus y)\oplus f(x\oplus z)$ for the same distribution. Both of these works used agreement theorems to reduce the biased case to the unbiased case: David et al. used the agreement theorem of Dinur and Steurer [DS14], and Dinur et al. proved their own agreement theorem, which we also use to prove Theorem 1.6.
@@
Filmus et al. [FLMM20], prompted by work of Nehama [Neh13] on judgment aggregation, proved Theorem 1.5 for the predicate $P_{\land}=\{(a,b,a\land b):a,b\in\{0,1\}\}$ and various distributions. Their analysis combined Bourgain’s tail bound [Bou02, KKO18] with a study of the “one-sided noise operator”.
Chase et al. [CFM+22] proved Theorem 1.5 for all predicates of the form $P_{f}=\{(a_{1},\dots,a_{m},f(a_{1},\dots,a_{m})):a_{1},\dots,a_{m}\in\{0,1\}\}$ for an arbitrary function $f\colon\{0,1\}^{m}\to\{0,1\}$, with respect to the uniform distribution. They asked whether Theorem 1.5 extends to arbitrary predicates, a question we answer in the affirmative in this paper. Their general approach was similar to the approach taken in this paper, combining Jones’ regularity lemma with the It Ain’t Over Till It’s Over theorem.
@@
We give an outline of our proof techniques in Section 2. We prove Theorem 1.7 in Section 3, Theorem 1.5 in Section 4, Theorem 1.6 in Section 5, Theorem 1.9 in Section 6, and Theorem 1.10 in Section 7. Our proofs require several versions of Jones’ regularity lemma [Jon16], proved in Section 8. We close the paper in Section 9 with a few open questions.
@@ 2.1 Monotone case
The proof of Theorem 1.7 has the same general outline as the proof of the triangle removal lemma [RS78] using Szémeredi’s regularity lemma [Sze78].
@@
The proof of Theorem 1.7 follows the same plan. Szémeredi’s regularity lemma is replaced by Jones’ regularity lemma [Jon16].
@@
Jones’ lemma is proved by a simple potential function argument. In applications we need a variant of this lemma for several functions, and also allowing an initial set as a starting point. We also need two extensions of the lemma, one to functions $f\colon\{0,1\}^{n}\to[0,1]$, for proving Theorem 1.9, and another to functions $f\colon\Sigma^{n}\to\Sigma$, for proving Theorem 1.10. Moreover, in order to obtain better parameters, in the full proof of Theorem 1.7 we use a version of Jones’ lemma which approximates $f$ by a decision tree rather than by a junta. We prove all of these different versions of Jones’ regularity lemma in Section 8.
@@
Applying the theorem to both $f$ and $1-f$, it implies that if $f$ is regular and its expectation is in $[\epsilon,1-\epsilon]$, then even if we sample all but a $q$-fraction of inputs, its expectation still lies in $[\delta,1-\delta]$ (where $\delta$ could be much smaller than $\epsilon$), with probability $1-\epsilon$. This explains its moniker.
@@ 2.2 General case
It turns out that the counting lemma does hold (under the additional assumption that $(\phi_{1},\dots,\phi_{m})$ is a $(\mu,\gamma)$-approximate generalized polymorphism) as long as $P$ satisfies no affine relations: there is no non-empty set $S$ such that $P|_{S}=\{x\in\{0,1\}^{S}:\bigoplus_{j\in S}x_{j}=b\}$. If $P$ satisfies the premise of Theorem 1.10 (for each $j\in[m]$ there is $x^{(j)}\in P$ which remains in $P$ after flipping the $j$’th coordinate) then this can be proved along the lines of the counting lemma for NAND. The general case requires an argument similar to the one used in [CFM+22], and also uses some ideas from [Mos12b].
In order to handle predicates with affine relations, we first prove Theorem 1.6, a generalization of linearity testing for arbitrary distributions. The proof uses the techniques of [DFH25]. Theorem 1.6 shows that if $j$ is a coordinate involved in an affine relation, then $f_{j}$ is close to a (possibly negated) XOR, and we fix $g_{j}$ to be this XOR. We then remove coordinates from $P$ until all affine relations disappear, enabling us to use the counting lemma.
@@ 3 Monotone case
In order to obtain the best possible parameters, we use Jones’ regularity lemma in the following form, essentially proved in [CFM+22]. For definiteness, we reprove it in Section 8.
@@
For every $\epsilon,\tau>0$, $m,d\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in(0,1)$, the following holds for $M=O(md/\epsilon\tau)$.
For all functions $f_{1},\dots,f_{m}\colon\{0,1\}^{n}\to\{0,1\}$ there exists a decision tree $T$ of depth at most $M$ such that for all $j$,
$\Pr_{\rho\sim T}[f_{j}|_{\rho}\text{ is $(d,\tau)$-regular with respect to $\mu_{p_{j}}$}]\geq 1-\epsilon,$
where $\rho$ is sampled by following $T$, sampling each variable encountered according to $\mu_{p_{j}}$.
@@ 3.1 Counting lemma
Let $q>0$ be a small enough parameter. We sample a random restriction $\rho\in(\{0,1,*\}^{m})^{n}$ by sampling each coordinate independently according to the following law, where $p_{j}=\Pr[\mu|_{j}=1]$, 00 is the zero vector, $e_{1}=(1,0,\dots,0)$, $s_{1}=(*,0,\dots,0)$, and $e_{j},s_{j}$ are defined analogously by making the special coordinate the $j$’th one:
@@
Given $\rho$, we can obtain a sample of $\mu$ by sampling a star in position $j$ according to $\mu|_{j}$. We denote this distribution by $\mu|\rho$.
@@ 3.2 Main result
Since no coordinate of $P$ is constant, we can apply Lemma 3.4 (the counting lemma) with $\epsilon:=\epsilon/2$ to obtain $d,\tau,\gamma$ for which the lemma holds. Note $d=O(\log(1/epsilon))$, $\tau=\Omega(\epsilon^{C})$, and $\delta=\Omega(\epsilon^{C})$.
We apply Theorem 3.3 (Jones’s regularity lemma) with $\epsilon:=\epsilon/2$, $p_{1},\dots,p_{m}$ given by $p_{j}=\Pr[\mu|_{j}=1]$, and the values of $d,\tau$ obtained from Lemma 3.4 to obtain a decision tree $T$ of depth $M=O(d/\epsilon\tau)=O((1/\epsilon)^{C})$.
We define the functions $g_{j}$ via the tree $T$ as follows. For each leaf $\rho\in T$ and $x\in\{0,1\}^{(\operatorname{dom}\rho)^{c}}$,
$g_{j}|_{\rho}(x)=\begin{cases}0&\text{if }f_{j}|_{\rho}\text{ is not $(d,\tau)$-regular or }\operatorname*{\mathbb{E}}_{\mu|_{j}}[f_{j}|_{\rho}]\leq\epsilon/2,\\ f_{j}|_{\rho}(x)&\text{otherwise}.\end{cases}$
If we sample $\rho$ according to $\mu|_{j}$ then according to Jones’ regularity lemma, $f_{j}|_{\rho}$ is $(d,\tau)$-regular (with respect to $\mu|_{j}$) with probability at least $21-\epsilon/$ 2. This shows that $\Pr_{\mu|_{j}}[g_{j}\neq f_{j}]\leq\epsilon/2+\epsilon/2=\epsilon$.
It remains to show that $(g_{1},\dots,g_{m})$ is an extended polymorphism of $P$ for small enough $\delta$. If this is not the case, then there exists a partial assignment $\rho\in T$ such that $(g_{1}|_{\rho|_{1}},\dots,g_{m}|_{\rho|_{m}})$ is not an extended polymorphism of $P$.
Apply Lemma 3.4 (the counting lemma) to $\phi_{j}=f_{j}|_{\rho|_{j}}$. Observe that for all $(y^{(1)},\dots,y^{(m)})\in\{0,1\}^{(\operatorname{dom}\rho)^{c}}$, $(g_{1}|_{\rho_{1}}(y^{(1)}),\dots,g_{m}|_{\rho_{m}}(y^{(m)}))\leq\alpha$, where $\alpha$ is the assignment defined in the lemma. Since $(g_{1}|_{\rho|_{1}},\dots,g_{m}|_{\rho|_{m}})$ is not an extended polymorphism of $P$ and $P$ is monotone, necessarily $\alpha\notin P$. Therefore the lemma shows that
$\Pr_{(y^{(1)},\dots,y^{(m)})\sim\mu^{(\operatorname{dom}\rho)^{c}}}[(f_{1}|_{\rho|_{1}}(y^{(1)}),\dots,f_{m}|_{\rho|_{m}}(y^{(m)}))\notin P]\geq\gamma,$
implying that $(f_{1},\dots,f_{m})$ is not a $(\mu,\delta)$-approximate generalized polymorphism for $\delta=\min(\mu)^{M}\gamma/2=1/\exp\Theta(1/\epsilon^{C})$, where $\min(\mu)=\min_{w\in\mu}(\mu(w))$. Choosing this value of $\delta$ completes the proof. ∎

### v2

@@ 1 Introduction
The classical BLR linearity test [BLR] states that if $f\colon\{0,1\}^{n}\to\{0,1\}$ satisfies $\Pr_{x,y\sim\mu_{1/2}}[f(x)\oplus f(y)=f(x\oplus y)]\geq 1-\epsilon$ then there exists a linear function $g\colon\{0,1\}^{n}\to\{0,1\}$ satisfying $\Pr_{x\sim\mu_{1/2}}[g(x)\neq f(x)]\leq\epsilon$; here $\mu_{1/2}$ is the uniform distribution over $\{0,1\}^{n}$. Linearity testing is a special case of a more general problem, approximate polymorphisms, introduced in [CFMMS2022], which unifies several other related results in the literature, including Kalai’s quantitative Arrow theorem [Kalai] and AND testing [FLMM2020].
@@
$\Pr_{\begin{subarray}{c}(x^{(1)}_{1},\dots,x^{(m)}_{1})\sim\mu\\ \cdots\\ (x^{(1)}_{n},\dots,x^{(m)}_{n})\sim\mu\end{subarray}}[(f(x^{(1)}),\dots,f(x^{(m)}))\in P]\geq 1-\epsilon.$
@@
Similarly, Keller’s quantitative Arrow theorem [Keller2012] (which improves on Kalai’s result) can be formulated as follows, in the special case of three candidates:
Theorem 1.3 (Quantitative Arrow theorem).
Let $\mathrm{NAE}_{3}=\{(a,b,c)\in\{0,1\}^{3}:(a,b,c)\neq(0,0,0),(1,1,1)\}$, and let $\mu$ be the uniform distribution over $\mathrm{NAE}_{3}$.
If $f\colon\{0,1\}^{n}\to\{0,1\}$ is a $(\mu,\epsilon)$-approximate polymorphism of $\mathrm{NAE}_{3}$ then there exists a polymorphism $g\colon\{0,1\}^{n}\to\{0,1\}$ of $\mathrm{NAE}_{3}$ such that $\Pr_{x\sim\mu_{1/2}}[g(x)\neq f(x)]=O(\epsilon^{1/3})$.
In this paper, we extend Theorems 1.2 and 1.3 to arbitrary predicates $P\subseteq\{0,1\}^{m}$ (for all $m$) and to arbitrary distributions $\mu$ on $P$ with full support.
One might hope for a result of the following form: Any approximate polymorphism of $P$ is close to a polymorphism of $P$. While this holds for some predicates $P$, it fails for others, as the following counterexample from [CFMMS2022] demonstrates.
@@ 1.1 Other results
Theorem 1.6 has an unspecified dependence between $\delta$ and $\epsilon$; the dependence arising from our current proof is of tower type. We can improve this dependence in the case of linearity testing.
@@
Another difference between Theorem 1.6 and Theorem 1.7 is that we obtain the additional guarantee that if $f_{i}=f_{j}$ and the corresponding marginals of $\mu$ coincide, then $g_{i}=g_{j}$. We are also able to obtain this guarantee in the case of monotone predicates. The guarantee implies that if all marginals of $\mu$ are identical then an approximate polymorphism is close to an exact polymorphism.
@@
Let $P\subseteq\{0,1\}^{m}$ be a non-empty monotone predicate: if $x\in P$ and $y\leq x$ (pointwise) then $y\in P$. Let $\mu$ be a distribution on $P$ with full support. For every $\epsilon>0$ there exists $\delta>0$ such that the following holds for all $n$.
@@
Fix $0<p<1/2$. For every $\epsilon>0$ there exist $\delta>0$ and $j\in\mathbb{N}$ such that the following holds for all $n$ such that $pn$ is an integer.
@@
We provide an alternative proof of Theorem 1.9 by applying Theorem 1.8 to the NAND predicate. Our proof (which uses the same reduction from $\binom{[n]}{pn}$ to the product measure $\mu_{p}$ on $\{0,1\}^{n}$) is arguably simpler.
@@ 1.2 Related work
Arrow’s celebrated theorem [Arrow] can be expressed in the language of polymorphisms. Let $m\geq 3$. For every permutation $\pi\in S_{m}$, let $I(\pi)\in\{0,1\}^{\binom{m}{2}}$ be the following vector: $I(\pi)(i,j)=[\pi(i)>\pi(j)]$. Arrow’s theorem states that if $f_{i,j}$ is a generalized polymorphism of $P_{m}:=\{I(\pi):\pi\in S_{m}\}$ and each $f_{i,j}$ is unanimous (satisfies $f_{i,j}(b,\dots,b)=b$ for $b\in\{0,1\}$) then there exists $k\in[n]$ such that $f_{i,j}(x)=x_{k}$ for all $i,j$.
Kalai [Kalai] considered the case $m=3$. He showed that if $f_{1,2},f_{2,3},f_{3,1}$ is a $(\mu,\epsilon)$-approximate polymorphism of $P_{3}$, where $\mu$ is the uniform distribution, and furthermore the $f_{i,j}$ are balanced (satisfy $\Pr_{\mu_{1/2}}[f_{i,j}=1]=1/2$), then there exists $k\in[n]$ such that either $f_{i,j}(x)=x_{k}$ for all $i,j$, or $f_{i,j}=1-x_{k}$ for all $i,j$. In this case the predicate $P_{3}$ consists of all triples $(a,b,c)$ such that $a,b,c$ are not all equal. Keller [Keller2010] extended Kalai’s result to arbitrary $m\geq 3$ and to arbitrary distributions, under various (severe) restrictions on the $f_{i,j}$.
Mossel [Mossel2012] determined all generalized polymorphisms of $P_{m}$ for all $m\geq 3$ (without assuming unanimity), and proved that approximate generalized polymorphisms of these predicates (with respect to the uniform distribution) are close to exact generalized polymorphisms. His techniques in fact work for arbitrary distributions (the missing piece, reverse hypercontractivity for arbitrary distributions, was proved in [MOS13]). Keller [Keller2012] improved on Mossel’s result in the case of the uniform distribution by determining the optimal dependence between $\epsilon$ and $\delta$.
@@
Blum, Luby and Rubinfeld [BLR] were the first to propose the BLR test. They analyzed it using self-correction. Bellare et al. [BCHKS96] later gave a different argument using Fourier analysis. The test was generalized to arbitrary prime fields by Kiwi [Kiwi03].
David et al. [DDGKS17] extended the BLR test to the setting of constant weight inputs, which is analogous to the setting of Theorem 1.9. Translated to the setting of Theorem 1.7, they extended the BLR test to a natural distribution $\mu$ with $\Pr[\mu|_{i}=1]=p$ for all $i$. Dinur et al. [DFH2025] analyzed the related “affine test” $f(x\oplus y\oplus z)=f(x)\oplus f(x\oplus y)\oplus f(x\oplus z)$ for the same distribution. Both of these works used agreement theorems to reduce the biased case to the unbiased case: David et al. used the agreement theorem of Dinur and Steurer [DS14], and Dinur et al. proved their own agreement theorem, which we also use to prove Theorem 1.7.
@@
Filmus et al. [FLMM2020], prompted by work of Nehama [Nehama] on judgment aggregation, proved Theorem 1.6 for the predicate $P_{\land}=\{(a,b,a\land b):a,b\in\{0,1\}\}$ and various distributions. Their analysis combined Bourgain’s tail bound [Bourgain02, KKO18] with a study of the “one-sided noise operator”.
Chase et al. [CFMMS2022] proved Theorem 1.6 for all predicates of the form $P_{f}=\{(a_{1},\dots,a_{m},f(a_{1},\dots,a_{m})):a_{1},\dots,a_{m}\in\{0,1\}\}$ for an arbitrary function $f\colon\{0,1\}^{m}\to\{0,1\}$, with respect to the uniform distribution. They asked whether Theorem 1.6 extends to arbitrary predicates, a question we answer in the affirmative in this paper. Their general approach was similar to the approach taken in this paper, combining Jones’ regularity lemma with the It Ain’t Over Till It’s Over theorem.
@@
We give an outline of our proof techniques in Section 2. We prove Theorem 1.8 in Section 3, Theorem 1.6 in Section 4, Theorem 1.7 in Section 5, Theorem 1.9 in Section 6, and Theorem 1.10 in Section 7. Our proofs require several versions of Jones’ regularity lemma [Jones2016], proved in Section 8. We close the paper in Section 9 with a few open questions.
@@ Acknowledgments
We thank Elchanan Mossel for pointing us to his unpublished work with Schramm [MosselSchramm2009].
We thank Yumou Fei for pointing out a mistake in the proof of Theorem 1.8. The original proof attempted to use a more efficient version of Jones’ regularity lemma, which outputs a decision tree rather than a junta (cf. [Mossel20]), in order to obtain better parameters.
@@ 2.1 Monotone case
The proof of Theorem 1.8 has the same general outline as the proof of the triangle removal lemma [RSz78] using Szémeredi’s regularity lemma [Sz78].
@@
The proof of Theorem 1.8 follows the same plan. Szémeredi’s regularity lemma is replaced by Jones’ regularity lemma [Jones2016]. [^1] The same regularity lemma had been proved earlier by Mossel and Schramm [MosselSchramm2009] and (independently) by O’Donnell, Servedio, Tang, and Wan [ODSTW2010], but Jones’ paper is the only one which is publicly available. Moseel [Mossel20] proved a version of the regularity lemma for several functions at once.
@@
Jones’ lemma is proved by a simple potential function argument. In applications we need a variant of this lemma for several functions, and also allowing an initial set as a starting point. We also need two extensions of the lemma, one to functions $f\colon\{0,1\}^{n}\to[0,1]$, for our alternative proof of Theorem 1.9, and another to functions $f\colon\Sigma^{n}\to\Sigma$, for proving Theorem 1.10. We prove all of these different versions of Jones’ regularity lemma in Section 8.
@@
Applying the theorem to both $f$ and 1f1-f, it implies that if $f$ is regular and its expectation is in $\epsilon,1-\epsilon]$, then even if we sample all but a $q$-fraction of inputs, its expectation still lies in $\delta,1-\delta]$ (where $\delta$ could be much smaller than $\epsilon$), with probability $1-\epsilon$. This explains its moniker.
@@ 2.2 General case
It turns out that the counting lemma does hold (under the additional assumption that $(\phi_{1},\dots,\phi_{m})$ is a $(\mu,\gamma)$-approximate generalized polymorphism) as long as $P$ satisfies no affine relations: there is no non-empty set $S$ such that $P|_{S}=\{x\in\{0,1\}^{S}:\bigoplus_{j\in S}x_{j}=b\}$. If $P$ satisfies the premise of Theorem 1.10 (for each $j\in[m]$ there is $x^{(j)}\in P$ which remains in $P$ after flipping the $j$’th coordinate) then this can be proved along the lines of the counting lemma for NAND. The general case requires an argument similar to the one used in [CFMMS2022], and also uses some ideas from [Mossel2012].
In order to handle predicates with affine relations, we first prove Theorem 1.7, a generalization of linearity testing for arbitrary distributions. The proof uses the techniques of [DFH2025]. Theorem 1.7 shows that if $j$ is a coordinate involved in an affine relation, then $f_{j}$ is close to a (possibly negated) XOR, and we fix $g_{j}$ to be this XOR. We then remove coordinates from $P$ until all affine relations disappear, enabling us to use the counting lemma.
@@ 3 Monotone case
In order to obtain the best possible parameters, we use Jones’ regularity lemma in the following form, essentially proved in [CFMMS2022]. For definiteness, we reprove it in Section 8.
@@
For every $\epsilon,\tau>0$, $m,d\in\mathbb{N}$ and $p_{1},\dots,p_{m}\in(0,1)$ there exists a parameter $M$ for which the following holds.
For all functions $f_{1},\dots,f_{m}\colon\{0,1\}^{n}\to\{0,1\}$ there exists a set $J$ of size at most $M$ such that for all $j$,
$\Pr_{x\sim\mu_{j}}[f_{j}|_{J\leftarrow x}\text{ is $(d,\tau)$-regular with respect to $\mu_{p_{j}}$}]\geq 1-\epsilon.$
@@ 3.1 Counting lemma
Let $q>0$ be a small enough parameter. We sample a random restriction $\rho\in(\{0,1,*\}^{m})^{n}$ by sampling each coordinate independently according to the following law, where $p_{j}=\Pr[\mu|_{j}=1]$, 0 is the zero vector, $e_{1}=(1,0,\dots,0)$, $s_{1}=(*,0,\dots,0)$, and $e_{j},s_{j}$ are defined analogously by making the special coordinate the $j$’th one:
@@
Given $\rho$, we can obtain a sample of $\mu$ by sampling a star in position $j$ according to $\mu|_{j}$. We denote this distribution by $\mu|_{\rho}$.
@@ 3.2 Main result
Since no coordinate of $P$ is constant, we can apply Lemma 3.4 (the counting lemma) with $\epsilon:=\epsilon/2$ to obtain $d,\tau,\gamma$ for which the lemma holds.
We apply Theorem 3.3 (Jones’s regularity lemma) with $\epsilon:=\epsilon/2$, $p_{1},\dots,p_{m}$ given by $p_{j}=\Pr[\mu|_{j}=1]$, and the values of $d,\tau$ obtained from Lemma 3.4 to obtain a set $J$ of size at most $M$.
We define the functions $g_{j}$ as follows. For each setting $x\in\{0,1\}^{J}$,
$g_{j}|_{J\leftarrow x}(y)=\begin{cases}0&\text{if }f_{j}|_{J\leftarrow x}\text{ is not $(d,\tau)$-regular or }\operatorname*{\mathbb{E}}_{\mu|_{j}}[f_{j}|_{J\leftarrow x}]\leq\epsilon/2,\\ f_{j}|_{J\leftarrow x}(y)&\text{otherwise}.\end{cases}$
If we sample $x$ according to $\mu|_{j}$ then according to Jones’ regularity lemma, $f_{j}|_{J\leftarrow x}$ is $(d,\tau)$-regular (with respect to $\mu|_{j}$) with probability at least $1-\epsilon/2$. This shows that $\Pr_{\mu|_{j}}[g_{j}\neq f_{j}]\leq\epsilon/2+\epsilon/2=\epsilon$.
It remains to show that $(g_{1},\dots,g_{m})$ is an extended polymorphism of $P$ for small enough $\delta$. If this is not the case, then there exists an assignment $(x^{(1)},\dots,x^{(m)})\in P^{J}$ such that $(g_{1}|_{J\leftarrow x^{(1)}},\dots,g_{m}|_{J\leftarrow x^{(m)}})$ is not an extended polymorphism of $P$.
Apply Lemma 3.4 (the counting lemma) to $\phi_{j}=f_{j}|_{J\leftarrow x^{(j)}}$. Observe that for all $(y^{(1)},\dots,y^{(m)})\in P^{J^{c}}$, $(g_{1}|_{J\leftarrow x^{(1)}}(y^{(1)}),\allowbreak\dots,\allowbreak g_{m}|_{J\leftarrow x^{(m)}}(y^{(m)}))\leq\alpha$, where $\alpha$ is the assignment defined in the lemma. Since $(g_{1}|_{J\leftarrow x^{(1)}},\allowbreak\dots,\allowbreak g_{m}|_{J\leftarrow x^{(m)}})$ is not an extended polymorphism of $P$ and $P$ is monotone, necessarily $\alpha\notin P$. Therefore the lemma shows that
$\Pr_{(y^{(1)},\dots,y^{(m)})\sim\mu^{J^{c}}}[(f_{1}|_{J\leftarrow x^{(1)}}(y^{(1)}),\dots,f_{m}|_{J\leftarrow x^{(m)}}(y^{(m)}))\notin P]\geq\gamma,$
implying that $(f_{1},\dots,f_{m})$ is not a $(\mu,\delta)$-approximate generalized polymorphism for $\delta=\min(\mu)^{M}\gamma/2$, where $\min(\mu)=\min_{w\in\mu}(\mu(w))$. Choosing this value of $\delta$ completes the proof. ∎